The length of a rectangle is two less than six times its width. The perimeter in 66 cm. Find the length and width of the rectangle. I am stuck. I keep getting here: 70=16w so w=4.375 and l=24.25. But when I do this it adds up to: 57.25 not 66
Answer by lwsshak3(11628)

The length of a rectangle is two less than six times its width. The perimeter in 66 cm. Find the length and width of the rectangle.
**
let W=width
let L=length
..
L=6W-2
2W+2L=perimeter
2W+2(6W-2)=66
2W+12W-4=66
14W=70
W=70/14
W=5
L=6W-2=
L=30-2
L=28
Check:
perimeter=2L+2W=56+10=66